OpenStudy (anonymous):

if ab+1/ab=4, what is the value of a^2b^2+1/a^2b^2?

zepdrix (zepdrix):

The question is very unclear. Did you mean \(\large\rm ab+\frac{1}{ab}=4\) or \(\large\rm \frac{ab+1}{ab}=4\) ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

(ab) + 1/ab = 4

zepdrix (zepdrix):

Hmm I remember seeing this question before D: Trying to rememberrrr

OpenStudy (anonymous):

I want to know the steps to solving it.

zepdrix (zepdrix):

Well let's try this :) What happens when you square each side?\[\large\rm \left(ab+\frac{1}{ab}\right)^2=4^2\]Do you remember how to expand out a pair of binomials like this? Sometimes we call it FOIL'ing.

OpenStudy (anonymous):

yes I know FOIL'ing

OpenStudy (anonymous):